# A Narrative Review of Identity, Data, and Location Privacy Techniques in Edge Computing and Mobile Crowdsourcing

### v2

@@ Abstract
As digital technology advances, the proliferation of connected devices poses significant challenges and opportunities in mobile crowdsourcing and edge computing. This narrative review focuses on the need for privacy protection in these fields, emphasizing the increasing importance of data security in a data-driven world. Through an analysis of contemporary academic literature, this review provides an understanding of the current trends and privacy concerns in mobile crowdsourcing and edge computing. We present insights and highlight advancements in privacy-preserving techniques, addressing identity, data, and location privacy. This survey serves as a useful resource for researchers, industry professionals, and policymakers, offering an overview of privacy challenges and potential solutions in these interconnected domains.
Keywords Privacy; Security; Edge Computing, Location, Mobile Crowd-sourcing, Cloud Computing
@@ 1 Introduction
The advent of the digital era has resulted in a significant surge in connected devices, generating a substantial volume of data transmitted over the Internet [1]. Managed predominantly via cloud infrastructures, the ubiquity of smart devices such as smartphones, tablets, smartwatches, and fitness trackers has woven them into the fabric of daily life. These devices gather vast amounts of contextual information about users, including their location, activities, and environmental conditions, crucial for applications designed to predict user behavior and provide personalized experiences [2, 3].
Mobile crowdsourcing has emerged as a significant phenomenon where individuals collectively contribute data through various digital platforms [4]. While applications such as traffic monitoring systems benefit from this crowdsourced data by offering real-time insights, the process raises substantial privacy concerns related to data transparency and the risks of unauthorized sensitive information access [5, 6].
Moreover, the exponential growth in mobile data traffic presents challenges in data management and processing [7]. Edge computing, which processes data closer to its source rather than relying solely on central servers, has been identified as a promising solution to alleviate network congestion. However, this decentralization also introduces potential privacy and security vulnerabilities that need to be addressed [8].
This paper explores the intersection of mobile crowdsourcing and edge computing with a particular focus on privacy preservation. We provide an overview of existing research addressing these technologies and their implications for privacy, aiming to equip researchers, practitioners, and policymakers with the necessary insights to navigate the complexities of ensuring privacy while advancing technological capabilities.
@@ 1.1 Literature Review Strategy
Our literature review strategy involved searches in academic repositories such as ACM Digital Library, IEEE Xplore, Springer, Elsevier, Google Scholar, DBLP, CiteSeerX, Microsoft Academic Search, Web of Science, and ScienceDirect. Publications from the period 2012 to 2021 were primarily considered, with a final selection of approximately 52 peer-reviewed papers from an initial pool of 150 papers based on relevance to mobile crowdsourcing and edge computing. These papers were analyzed for trends, methodologies, findings, and gaps in research.
Search Queries: Structured queries were employed to search for relevant literature, focusing on privacy preservation, data security, and the integration of mobile crowdsourcing with edge computing, among other aspects.
1. ("Mobile Crowdsourcing" AND "Edge Computing" AND "Privacy Preservation") AND ("Data Security" OR "Real-Time Data Processing") AND (2012-2021)
@@
3. ("Mobile Crowdsourcing" AND "Edge Computing") AND ("Privacy" AND "Information Overload") AND ("Smart Devices" OR "User Behavior Prediction") AND (2012-2021)
@@ 1.2 Contributions
This work contributes to the field:
@@
    It offers an interdisciplinary examination of mobile crowdsourcing and edge computing, emphasizing the importance of privacy preservation.
@@
    It synthesizes both contemporary and foundational literature, providing an understanding of identity privacy, data privacy, and location privacy (based on their relevance to importance) in these interconnected domains, and identifying current research trends.
A meta-review of previous surveys with focus on privacy in these individual domains is presented in Table 1.
@@
Edge Computing Real-time and local data processing, IoT integration, Edge AI, Security Studies highlighted capabilities for local processing, integration with AI, and identified gaps in privacy protection specific to edge computing environments. Key studies include [9], [10], [11], [12], [13], [8], [14], [15].
@@
Mobile Computing Security Trends, Healthcare, Cloud Integration, AR Applications Focused on evolving security challenges, particularly in mobile cloud contexts and healthcare applications during the COVID-19 pandemic, and AR enhancements. Relevant studies include [16], [17], [18], [19], [20].
1.3 Privacy Concerns in Mobile Crowdsourcing and Edge Computing
@@
Next, we discuss mobile crowdsourcing and edge computing and examine the privacy challenges in these domains, focusing on identity privacy, data privacy, and location privacy due to their significant impact on protecting users in these interconnected systems.
Identity Privacy
@@
Identity privacy refers to the protection of an individual’s identity from being exposed or inferred [21] during the process of mobile crowdsourcing or edge computing activities. This includes safeguarding personal information such as names, social security numbers, and other identifiers that can link activities to a specific individual.
@@
Mobile crowdsourcing often involves collecting data from numerous participants. Ensuring that these participants remain anonymous or pseudonymous is crucial to prevent misuse of personal information [22]. Exposing personal identifiers can lead to identity theft, where malicious actors could use the stolen identities to commit fraud or other illegal activities [23]. Protecting identity privacy is fundamental in establishing and maintaining trust among users.
Data Privacy
Data privacy refers to the protection of the actual data collected, processed, and stored [24] during mobile crowdsourcing and edge computing. This involves ensuring that sensitive information remains confidential and is not accessed, altered, or disclosed without authorization.
Crowdsourced data can include sensitive information, such as health data, financial transactions, or personal communications. Protecting this data from unauthorized access and ensuring its integrity is highly important [24]. Various regulations, such as the General Data Protection Regulation (GDPR) or the Health Insurance Portability and Accountability Act (HIPAA), mandate strict data privacy measures. Data breaches can have severe consequences, including financial loss, reputational damage, and loss of user trust.
Location Privacy
Location privacy involves the protection of information regarding an individual’s geographical location obtained through mobile crowdsourcing and edge computing. This includes ensuring that location data cannot be used to track or monitor an individual’s movements without their consent.
@@
Unauthorized access to location data can lead to stalking, harassment, or other malicious activities where an individual’s movements are tracked without their knowledge or consent [25]. Disclosure of sensitive locations, such as homes, workplaces, or places of interest, can pose security risks and privacy violations [26]. Users should have control over who can access their location data and for what purposes. It is, therefore, important to ensure the location privacy for the sake if users’ autonomy and their right to control their personal information [27].
@@ 2 Mobile Crowdsourcing
Mobile crowdsourcing represents a field where collective user contributions, primarily via mobile devices, drive data collection and service provision [28, 29]. By leveraging the ubiquitous nature of mobile devices, this domain taps into a vast pool of user-submitted data, encompassing geospatial and behavioral data as well as environmental and health metrics. Its applications are vast, spanning environmental monitoring, disaster response, urban planning, healthcare, and transportation, highlighting the versatility and impact of mobile crowdsourcing [30].
@@
Figure 1: Anatomy of a mobile crowdsourcing campaign
Key Stages in Mobile Crowdsourcing: We show the key stages in mobile crowdsourcing in Figure 1 .
@@
    Tasking Stage: This initial phase involves defining clear and actionable data collection objectives that align with the overarching goals of the crowdsourcing campaign.
@@
    Collecting Stage: At this juncture, data is actively acquired from users through their mobile devices. This data can range from environmental readings to user-generated content and location data, providing a rich dataset for analysis.
@@
    Storing Stage: The collected data is centralized on cloud or edge servers for processing. This stage is crucial for managing the volume of data and ensuring it is accessible for analysis.
@@
    Mining Stage: Here, advanced analytics and data mining techniques are applied to the aggregated data to extract meaningful insights, patterns, and trends.
@@
Technological Frameworks and Challenges: The technological infrastructure supporting mobile crowdsourcing includes a range of components from mobile devices and network communication protocols to cloud and edge computing platforms. These technologies facilitate the efficient collection, transmission, and processing of crowdsourced data. However, this complex ecosystem also introduces challenges related to scalability, data quality control, and the integration of heterogeneous data sources.
Privacy Considerations: The expansive data collection inherent to mobile crowdsourcing initiatives raises significant privacy concerns. The transmission of sensitive personal information, such as location data and personal identifiers, to centralized servers poses risks related to data security and user privacy [32]. Addressing these concerns, the field is increasingly looking towards edge computing as a viable solution. By processing data closer to its source, edge computing offers enhanced performance, reduced transmission costs, and improved data security, presenting a promising avenue for mitigating privacy risks in mobile crowdsourcing [33].
Literature: In mobile computing, recent surveys have focused on security trends, healthcare applications during the COVID-19 pandemic, and the integration with cloud computing. A study [16] explores evolving security challenges in mobile cloud computing and strategies for safeguarding these environments. The significant role of mobile computing in healthcare, especially during the pandemic, is highlighted in a survey [17], addressing both challenges and potential solutions. Additionally, the integration of mobile computing with edge computing is examined in [18], where the authors discuss the benefits of leveraging edge computing for reducing latency and improving processing capabilities in mobile applications. The survey [19] delves into the advancements in mobile augmented reality (AR), focusing on the integration of mobile computing with AR to enhance user experiences and application performance. Further research [20] reviews the application of mobile computing in smart cities.
Challenges: As mobile crowdsourcing continues to evolve, future research and development efforts are likely to focus on enhancing privacy protections, improving data processing efficiencies, and expanding the applicability of crowdsourcing models. Innovations in encryption, decentralized data processing, and user-centric privacy controls are expected to play pivotal roles in shaping the next generation of mobile crowdsourcing solutions. Through the collective power of mobile device users, mobile crowdsourcing can bring transformative changes across numerous domains. However, ensuring the ethical and secure handling of user-contributed data remains a critical challenge that must be addressed to realize the full potential of this innovative approach.
@@ 2.1.1 Identity Privacy in Mobile Crowdsourcing
Identity privacy in mobile crowdsourcing is a crucial aspect of safeguarding user data, which includes sensitive information like names, addresses, and contact details. Discussions on identity privacy focus on the implementation of robust forward security measures to protect cloud data [34, 22]. Identity-based encryption methods are highlighted as a means to secure cloud data, showcasing the multifaceted approach required to enhance privacy protection in cloud computing contexts [23]. The Consolidated Identity Management Systems (CIDM) are proposed to safeguard user identity by separating authorization credentials, adding an extra authentication layer, and securing communication links with cloud service providers [35]. Further discussions emphasize blockchain-based trust models and strategies for identity management that negate the need for trusted third parties [36, 37].
2.2 Data Privacy in Mobile Crowdsourcing
@@
2.3 Location Privacy in Mobile Crowdsourcing
@@
Collectively, these research efforts illustrate the diverse strategies being explored to protect identity, data, and location privacy in cloud computing. By leveraging encryption, spatial cloaking, and cryptographic authentication, these methods offer promising avenues for securing sensitive information against unauthorized access and misuse, thereby contributing to the broader goals of privacy and security in the digital age.
These methods are summarized in Table 2 .
Table 2: Privacy Concerns and Solutions in Mobile Cloud Computing

### v3

# A Narrative Review of Identity, Data and Location Privacy Techniques in Edge Computing and Mobile Crowdsourcing

@@ Abstract
As digital technology advances, the proliferation of connected devices poses significant challenges and opportunities in mobile crowdsourcing and edge computing. This narrative review focuses on the need for privacy protection in these fields, emphasizing the increasing importance of data security in a data-driven world. Through an analysis of contemporary academic literature, this review provides an understanding of the current trends and privacy concerns in mobile crowdsourcing and edge computing. We present insights and highlight advancements in privacy-preserving techniques, addressing identity, data, and location privacy. This review also discusses the potential directions that can be useful resources for researchers, industry professionals, and policymakers.
@@ 1 Introduction
The advent of the digital era and information overload has resulted in a significant increase in connected devices, generating a substantial volume of data transmitted over the Internet [8]. Managing predominantly through cloud infrastructures, the ubiquity of smart devices such as smartphones, tablets, smartwatches, and fitness trackers has made it even more complex. These devices routinely gather extensive contextual information about users, including their locations, activities, and environmental conditions—data that are vital for applications aimed at predicting user behavior and delivering personalized experiences [4, 49].
In response to these developments, mobile crowdsourcing has emerged as a pivotal solution. This approach involves individuals collectively contributing data via various digital platforms [36]. Applications such as traffic monitoring systems leverage crowdsourced data to provide real-time insights. However, these methods also raise significant concerns regarding data privacy and the risks of unauthorized access to sensitive information [7, 19].
The rapid increase in mobile data traffic poses substantial challenges for data management and processing. Edge computing has been identified as a promising solution to these challenges by processing data closer to its source rather than relying solely on central servers. This approach helps alleviate network congestion but also introduces new privacy and security vulnerabilities that must be addressed [67].
This paper explores the intersection of mobile crowdsourcing and edge computing with a particular emphasis on privacy preservation. It highlights the potential for integrating these technologies to enhance data processing efficiency while maintaining user privacy. We provide an overview of existing research on these technologies and their implications for privacy. Our goal is to equip researchers, practitioners, and policymakers with the necessary insights to navigate the complexities of ensuring privacy while advancing technological capabilities. Additionally, we discuss potential strategies for mitigating privacy risks associated with decentralized data processing and propose future research directions to address these challenges comprehensively.
@@ 1.1 Literature Review Strategy
Our literature review strategy involved searches in academic repositories such as ACM Digital Library, IEEE Xplore, Springer, Elsevier, Google Scholar, DBLP, CiteSeerX, Microsoft Academic Search, Web of Science, and ScienceDirect. Publications from the period 2012 to 2021 were primarily considered, with a final selection of approximately 52 peer-reviewed papers from an initial pool of 150 papers based on relevance to mobile crowdsourcing and edge computing. These papers were analyzed for trends, methodologies, findings, and gaps in research. The publication trend in the last few years is given in Figures 13.
Figure 1: Publication trend on privacy in edge computing and mobile crowdsourcing in recent years. The full year 2024 is not covered, so additional publications are expected.
Figure 2: Top publishers by number of publications in the field.
Figure 3: Publication trends over time by topic.
We used queries to search for relevant literature , shown in Listing 1:
Listing 1: Search Query Strings for Mobile Crowdsourcing and Edge Computing Studies
1. ("Mobile Crowdsourcing" AND "Edge Computing" AND "Privacy Preservation") AND ("Data Security" OR "Real-Time Data Processing") AND (2012-2023)
@@
3. ("Mobile Crowdsourcing" AND "Edge Computing") AND ("Privacy" AND "Information Overload") AND ("Smart Devices" OR "User Behavior Prediction") AND (2012-2023)
@@ 1.2 Contributions
This work contributes the following to the field:
@@
    This review offers an overview of mobile crowdsourcing and edge computing, with an emphasis on the importance of privacy preservation.
@@
    We synthesize both contemporary and foundational literature to provide an understanding of identity privacy, data privacy, and location privacy (based on their relevance to importance) and identify current research trends.
A meta-review of previous surveys is presented in Table 1, and the main keywords used in this work are given in Table 4.
@@
Edge Computing Real-time and local data processing, IoT integration, Edge AI, Security Studies highlighted capabilities for local processing, integration with AI, and identified gaps in privacy protection specific to edge computing environments. Key studies include [6, 38, 53, 42, 16, 67, 63, 66].
@@
Mobile Computing Security Trends, Healthcare, Cloud Integration, AR Applications Focused on evolving security challenges, particularly in mobile cloud contexts and healthcare applications during the COVID-19 pandemic, and AR enhancements. Relevant studies include [51, 5, 22, 20, 34].
1.3 Necessity of This Review
This review addresses a literature gap that surrounds the rapidly evolving fields of mobile crowdsourcing and edge computing. As our reliance on data-driven systems continues to grow, this review provides a timely analysis of current trends, privacy concerns, and research gaps in these interconnected domains. We also shed some light on key privacy issues, including those related to identity, data, and location, and offer some information to researchers, industry professionals, and policymakers. The flow of this review is shown in Figure 4.
Figure 4: Flow of the privacy-preserving methods covered in this work.
2 Privacy Concerns in Mobile Crowdsourcing and Edge Computing
@@
We discuss mobile crowdsourcing and edge computing and examine privacy challenges in these domains, focusing on identity privacy, data privacy, and location privacy due to their significant impact on protecting users in these interconnected systems.
2.1 Identity Privacy
@@
Identity privacy is crucial in protecting individual personal information from being exposed or inferred during mobile crowdsourcing or edge computing activities [64]. This protection encompasses safeguarding personal details such as names, social security numbers, and other identifiers that could link activities to a specific individual.
@@
In mobile crowdsourcing, data are often collected from numerous participants, making it essential to ensure their anonymity or pseudonymity. This is critical to prevent the misuse of personal information [56]. If personal identifiers are exposed, it can lead to identity theft, where malicious actors might use stolen identities to commit fraud or other illegal activities [54]. Thus, safeguarding identity privacy is fundamental for establishing and maintaining user trust.
The complexity of data collection processes in mobile crowdsourcing also raises concerns about implied consent and the need for more transparent consent management. Often, users agree to data collection practices without fully understanding the implications, inadvertently compromising their privacy. Therefore, it is vital to develop clearer consent mechanisms that ensure users are informed and their privacy is respected throughout the data collection process.
2.2 Data Privacy
Data privacy in the context of mobile crowdsourcing and edge computing involves protecting the data collected, processed, and stored from unauthorized access, alteration, or disclosure [58]. This protection is crucial, as crowdsourced data often include sensitive information such as health records, financial transactions, and personal communications. Ensuring the confidentiality and integrity of such data is vital to prevent unauthorized access and maintain user trust [58].
Various regulations, such as the General Data Protection Regulation (GDPR) (https://gdpr-info.eu/) (accessed on 22nd October, 2024) and the Health Insurance Portability and Accountability Act (HIPAA) (https://www.hhs.gov/hipaa/) (accessed on 22nd October, 2024), mandate strict measures to safeguard data privacy. Breaches can lead to severe consequences, including financial losses, reputational damage, and erosion of user trust. A significant concern in this domain is the risk associated with third-party data sharing. Once data enter the crowdsourcing ecosystem, users often lose control over their personal information as it flows through multiple parties in edge computing environments. This lack of control raises privacy risks that need to be addressed.
Furthermore, mobile crowdsourcing systems are vulnerable to specific attacks like Sybil attacks and data poisoning [30]. These attacks can manipulate the crowdsourcing process and potentially expose sensitive user information. Addressing these vulnerabilities requires robust security measures to protect against unauthorized data access and ensure data integrity throughout the crowdsourcing lifecycle.
2.3 Location Privacy
@@
Location privacy is the protection of information regarding an individual geographical location, particularly as it is obtained through mobile crowdsourcing and edge computing. This involves ensuring that location data cannot be used to track or monitor an individual’s movements without their explicit consent. Unauthorized access to location data poses significant risks, including stalking, harassment, and other malicious activities where an individual’s movements might be tracked without their knowledge or consent [35]. The disclosure of sensitive locations, such as homes, workplaces, or places of interest, can lead to security risks and privacy violations [69]. Therefore, it is crucial for users to have control over who can access their location data and for what purposes. Ensuring location privacy is fundamental to preserving users’ autonomy and their right to control their personal information [28].
The risks associated with sharing location data are many, for example, malicious actors can use precise location information to deduce behavioral patterns and predict where someone might be at any given time, potentially facilitating physical tracking or blackmail [35]. Additionally, unauthorized sharing or breaches of location data can lead to unintended surveillance and discrimination by authorities or third parties [62].
To mitigate these risks, it is essential for organizations to implement robust privacy measures. These include obtaining explicit user consent before enabling location-tracking features and ensuring transparency about how location data are used and shared. Techniques such as spatial cloaking and cryptographic methods can also enhance location privacy by anonymizing data and securing communication channels [14].
Adhoc Solutions
To address privacy concerns in mobile crowdsourcing and edge computing, researchers and practitioners are developing several innovative solutions.
One approach is the implementation of decentralized systems [13], which enhance user control over data and minimize the risks associated with centralized data repositories. This helps prevent vulnerabilities that arise from having a single point of failure or attack. Organizations are also shifting towards using advanced encryption techniques [11] to protect data both when it is stored and during transfer. The goal is to ensure that even if someone intercepts the data, it remains unintelligible without proper authorization, thereby safeguarding sensitive information. Technologies like differential privacy [23] help in protecting sensitive information while still allowing for meaningful data analysis. By adding noise to datasets, these techniques make it difficult to identify individual users, thus preserving privacy while enabling valuable insights.
There is an increasing focus on creating clear and easy-to-use systems for managing consent [70]. These frameworks ensure that users are fully informed about data-sharing practices and can easily control their preferences. This empowers users to have more say over their personal information and its handling, especially on crowdsourcing platforms. Also, leveraging blockchain technology [30] provides traceability and immutability in data handling practices. Blockchain enhances security and transparency, allowing users to verify how their data are used and shared. This not only improves security but also builds trust by providing a transparent record of transactions and data usage.
@@ 3 Mobile Crowdsourcing
Mobile crowdsourcing represents a field where collective user contributions, primarily via mobile devices, drive data collection and service provision [17, 50]. Mobile crowdsourcing leverages the ubiquitous nature of mobile devices into a vast pool of user-submitted data, encompassing geospatial and behavioral data as well as environmental and health metrics. Its applications are vast, covering environmental monitoring, disaster response, urban planning, healthcare, and transportation, and highlight the versatility and impact of mobile crowdsourcing [29].
@@
We show the key stages in mobile crowdsourcing in Figure 5.
@@
    Tasking Stage: This initial phase involves defining clear and actionable data collection objectives that align with the overarching goals of crowdsourcing.
@@
    Collecting Stage: At this point, the data are actively acquired from users through their mobile devices. The data can range from environmental readings to user-generated content and location data, providing a rich dataset for analysis.
@@
    Storing Stage: The collected data are centralized on cloud or edge servers for processing. This stage is crucial for managing the volume of data and ensuring it is accessible for analysis.
@@
    Mining Stage: Advanced analytics, machine learning (ML), and data mining techniques are applied to the aggregated data to extract meaningful insights, patterns, and trends.
@@
The technological infrastructure supporting mobile crowdsourcing includes a range of components from mobile devices and network communication protocols to cloud and edge computing platforms. These technologies facilitate the efficient collection, transmission, and processing of crowdsourced data. However, this complex ecosystem also introduces challenges related to scalability, data quality control, and the integration of heterogeneous data sources. The transmission of sensitive personal information, such as location data and personal identifiers, to centralized servers poses risks related to data security and user privacy [2, 3].
Figure 5: Anatomy of a mobile crowdsourcing campaign.
Literature: In mobile computing, recent surveys have focused on security trends, healthcare applications during the COVID-19 pandemic, and the integration with cloud computing. One study [51] explored evolving security challenges in mobile cloud computing and strategies for safeguarding these environments. The significant role of mobile computing in healthcare, especially during the pandemic, was highlighted in a survey [5], addressing both challenges and potential solutions. Additionally, the integration of mobile computing with edge computing is examined in [22], where the authors discuss the benefits of leveraging edge computing for reducing latency and improving processing capabilities in mobile applications. The survey [20] delves into the advancements in mobile augmented reality (AR), focusing on the integration of mobile computing with AR to enhance user experiences and application performance. Further research [34] reviews the application of mobile computing in smart cities.
@@ 3.1.1 Identity Privacy in Mobile Crowdsourcing
Identity privacy in mobile crowdsourcing is an important aspect of safeguarding user data, which includes sensitive information like names, addresses, and contact details. Discussions on identity privacy focus on the implementation of robust forward security measures to protect cloud data [31, 56]. Identity-based encryption methods are highlighted as a means to secure cloud data, which shows the multifaceted approach required to enhance privacy protection in cloud computing contexts [54]. Consolidated Identity Management Systems (CIDM) are proposed to safeguard user identity by separating authorization credentials, adding an extra authentication layer, and securing communication links with cloud service providers [27]. Further discussions emphasize blockchain-based trust models and strategies for identity management that negate the need for trusted third parties [26, 10].
3.1.2 Data Privacy in Mobile Crowdsourcing
@@
3.1.3 Location Privacy in Mobile Crowdsourcing
@@
Collectively, these research efforts illustrate the diverse strategies being explored to protect identity, data, and location privacy in cloud computing. These methods are summarized in Table 2.
Table 2: Privacy concerns and solutions in mobile cloud computing.